OpenStudy (anonymous):

What is the maximum height to which the ball rises if a spring gun projects a golf ball at an angle of 45 above the horizontal? The horizontal range is 10m.

OpenStudy (anonymous):

sir can you help me with this? given is kinda lack, i don't know what formula to use

OpenStudy (anonymous):

you should divide the velocity to 2 components, one of them is vertical the other is horizontal. For vertical component, Vx= V*Cos(45) you can find the vertical component. After that, you know the range, so you should multiply Vx with time t. then you get V*t you can use this for vertical calculation...

OpenStudy (anonymous):

still kinda dizzy so i get Vox=V*cos45 and Voy=V*sin45 what will i do with this?.. x.x

OpenStudy (anonymous):

x = v cos 45 . t 10 = ½√2 vt then, you can use this formulas vy = voy - gt vy² = voy² - 2gh with assume that vy = 0 at maximum height

OpenStudy (anonymous):

wait ill try to solve.. ill just ask if something bothers me thanks for the help guys :)

OpenStudy (anonymous):

@surygita the formula \[10=\frac{ 1 }{ 2 }\sqrt{2}vt\] is it from the formula \[h=vot \pm \frac{ 1 }{ 2 } g t ^{2}\] ?

OpenStudy (anonymous):

woops, I made my mistake